OpenStudy (anonymous):

PLEASE HELP!! A coin is tossed and a single 6-sided die is rolled. Find the probability of landing on the tail side of the coin and rolling a 5 on the die.

OpenStudy (thomas5267):

What is the probability of a coin landing on the tail side and what is the probability of rolling a 5 on a die?

OpenStudy (anonymous):

I think you have to divide 5 to 6

OpenStudy (anonymous):

is it 1/2

OpenStudy (thomas5267):

So you reckon that there is a 5-out-of-6 chance that the dice will land on 5?

OpenStudy (anonymous):

I don't really now but im sure its 1/2

OpenStudy (_michaeld22):

For the coin, there are 2 sides. The probability of landing on tails is 1/2, or 50%. There is only one '5' on a die. Six sides. So the probability of rolling a 5 is 1/6.

OpenStudy (anonymous):

Probability of head and 3 on a die is: P( head and 3 ) = p(head )* P(3) We multiply probabilities because we have 2 independent events. We know that p(head) = 1/2 Also p(3) = 1/6 Then P( head and 3) = 1/2 * 1/6 = 1/12